A large college class is graded on a total points system. The total points earned in a semester by the students in the class vary normally with a mean of 675 and a standard deviation of 50. Another large class in a different department is graded on a 0 to 100 scale. The final grades in that class follow a normal model with a mean of 82 and a standard deviation of 6. Aseda earns 729 points in the first class, while Ayeyi scores 90 in the second class. Which student did better and why?
1
Expert's answer
2020-08-25T12:58:09-0400

The criterion of better student performance in this case is deviation from the mean. The farther is student from the mean, the more outstanding he/she is. Let's calculate how far from the mean is each of students in terms of deviation.

For Aseda: "\frac{729-675}{50}=1.08". So, "\sigma_1 = 1.08 \sigma_0" .

For Ayeyi: "\frac{90-82}{6}=1.33". So, "\sigma_2 = 1.33 \sigma_0" .

Answer: Ayeyi did better, because her grade is "1.33 \sigma_0" away from the mean. Aseda did worse, because her grade is only "1.08 \sigma_0" away from the mean.